# Effects of lower floating-point precision on scale-resolving numerical simulations of turbulence

### v1

@@ Abstract
Modern computing clusters offer specialized hardware for reduced-precision arithmetic that can speed up the time to solution significantly. This is possible due to a decrease in data movement, as well as the ability to perform arithmetic operations at a faster rate. However, for high-fidelity simulations of turbulence, such as direct and large-eddy simulation, the impact of reduced precision on the computed solution and the resulting uncertainty across flow solvers and different flow cases have not been explored in detail and limits the optimal utilization of new high-performance computing systems. In this work, the effect of reduced precision is studied using four diverse computational fluid dynamics (CFD) solvers (two incompressible, Neko and Simson, and two compressible, PadeLibs and SSDC) using four test cases: turbulent channel flow at $Re_{\tau}=550$ and higher, forced transition in a channel, flow over a cylinder at $Re_{D}=3900$, and compressible flow over a wing section at $Re_{c}=50000$. We observe that the flow physics are remarkably robust with respect to reduction in lower floating-point precision, and that often other forms of uncertainty, due to for example time averaging, often have a much larger impact on the computed result. Our results indicate that different terms in the Navier–Stokes equations can be computed to a lower floating-point accuracy without affecting the results. In particular, standard IEEE single precision can be used effectively for the entirety of the simulation, showing no significant discrepancies from double-precision results across the solvers and cases considered. Potential pitfalls are also discussed.
@@ 1 Introduction
Computational fluid dynamics (CFD) has become an essential tool in both academic research and industry, spanning a diverse array of applications. Over time, a variety of models and numerical methods have been developed to integrate the governing equations of fluid motion, each tailored to specific use cases, desired accuracy, and computational constraints. Yet, until rather recently, all of them typically relied on IEEE double precision floating point numbers (FP64) to numerically compute the solution.
Due to recent shifts in hardware manufacturing tailored to decrease the energy consumption of floating-point operations and drive up performance for artificial intelligence (AI) applications, hardware support for lower precision floating-point numbers has become increasingly prevalent [1]. On these new platforms, lower precision offers both higher performance, higher energy efficiency, as well as a smaller memory footprint, reducing the amount of data movement necessary. All of these developments in the end lead to significant savings in both time and energy and thus provide large monetary savings for large-scale computations. As CFD practitioners, where most codes are limited by memory bandwidth due to large relatively sparse linear algebra systems, the use of FP32 instead of FP64 would, for example, move the bandwidth roofline by a factor 2, indicating an up to 2X of performance would be attainable for a bandwidth limited code [2]. A simple roofline comparison between two relatively common GPUs, Nvidia A100 and Nvidia GeForce RTX4080, is shown in Figure 1. The simple performance model clearly illustrates how the roofline for the RTX4080, with a significantly lower performance for FP64, improves for FP32 and is comparable to the more expensive A100. Considering that the operational intensity $I$ for a given code also improves with lower precision, this points to the significant gains that can be enabled by lower precision. In practice, a factor of 2×2$\times$2 × in performance is rarely achieved due to other factors such as kernel launch latencies or communication overhead in parallel communications.
@@
While there are clear opportunities for performance gains, an important question is to what extent these lower floating point formats can be used for scale-resolving simulation of turbulence without sacrificing the required accuracy. This is what we aim to assess in this work.
@@
This question is under active investigation by the research community and can be viewed as part of a general trend across many disciplines [3]. A full review of works on lower-precision arithmetics and CFD is outside the scope of this article, but a few selected references are provided for the benefit of the reader. The first example is from simulations of weather and climate. Several articles from Klöwer, Paxton, Düben, and Palmer et al. [4, 5, 6] highlight that lower precision and less exact computing open up for large performance improvements. In the realm of compressible flows, the CFD solver PyFR, [7, 8], used single precision (FP32) in the context of a high-order flux-reconstruction scheme for implicit large eddy simulation for a wide range of cases. Methods that enable mixed-precision computing while maintaining acceptable levels of accuracy have been proposed and investigated, for instance, in the context of the finite volume solver OpenFOAM [9], further encouraging results using OpenFOAM were also found by Brogi et al. [10]. Haridas et al. [11] explored the possibility of using neural networks to correct errors introduced using reduced precision arithmetic in simulating fluid dynamic problems. The feasibility of conducting mixed-precision operations in the context of high-order compact schemes were investigated by Song et al. [12] using the PadeLibs code for CFD. Wang et al. [13] proposed a mixed precision strategy in the finite volume method for unstructured grids that used high precision near solid bodies and lower precision far away from them. Freytag et al. [14] studied the performance and power efficiency of using reduced- and mixed-precision arithmetic for CFD. Bhola and Duraisamy [15, 16] performed analysis of errors incurred due to rounding in mixed precision computations. Walden et al. [17] investigated the speedup obtained on GPUs with the FUN3D CFD code by using reduced-precision arithmetic for their memory-bound linear solver kernel. Grout [18] used reduced-precision for the low-order time integration methods that were used to build higher-order methods in the so-called spectral differed correction method used in combustion CFD and studied the resulting rate of convergence. In order to reduce the communication bottleneck, Rogowski et al. [19] investigated reduced precision for communication in the context of Discontinuous Galerkin method using the SSDC CFD code. Finally, for CFD based on the Lattice Boltzmann method rather than solving the Navier-Stokes (NS) equations, [20] demonstrates the possibility to accelerate the solver using lower precision. It is also worth noting that finite volume codes used in industry, both open-source and commercial, offered the possibility to compile in FP32 long before the recent increase in low-precision computer hardware. There are also many mixed-precision algorithms for important linear algebra operations used in CFD, such as preconditioners and iterative solvers, to obtain a result in full double precision [21, 22, 3, 23]. This potentially suggests a wider scale of adoption of FP32 than what is documented in the scientific literature.
@@
In spite of these success stories, it is important to note that due to the nonlinear nature of the NS equations, cases where numerical precision will have a larger-than-expected impact can be found. In particular, in the context of dynamical systems, if there exist several attractors of the flow or a symmetry that is sensitive to small disturbances, lower floating-point precision can be detrimental to the validity of the simulation [24, 25]. However, if the simulation only has one attractor, and the results are insensitive to small thermal fluctuations, which in many ways can act similarly to low-precision [26], lower precision might be utilized effectively.
@@
The question in the focus of this article is whether reduced precision arithmetic can be successfully employed in direct numerical simulations (DNS) and wall-resolved large-eddy simulation (LES). The impression of the authors is that the DNS community is generally skeptical towards abandoning FP64. There are good reasons behind that. DNS are usually conducted to obtain the ground truth for a given flow and study all possible subtleties in its behavior. Therefore, even small errors are considered unacceptable. Furthermore, a new DNS is usually conducted at the limit of the computing budget, and simply trying out a low-precision run can be perceived as not worth the risk.
@@
Nevertheless, the main outcome of the DNS is often primarily expressed in a statistical description of the flow, and an argument can be made that it must, to a degree, be robust to small numerical errors, the origin of which may be both discretization and the precision of arithmetics. Indeed, there are already documented successful attempts in using FP32 for DNS of homogeneous isotropic turbulence (HIT) [27, 28, 29, 30], with no discernible differences observed between the FP32 and FP64 runs. While this is very encouraging, to trigger a shift across the wider high-fidelity CFD community, it is necessary for similar evidence to emerge across a broader range of flow cases and numerical techniques. The goal of the current work is to be a step in that direction, and toward a deeper understanding of the potential impact of lower floating precision on high-fidelity CFD. This builds on some preliminary work performed by Karp et al. [31, 32, 33].
@@
To that end, our study makes use of four different solvers (Neko, SSDC, Simson, PadeLibs), that differ in both the formulation of the governing equations and the approach to discretization. We consider four use cases, with a focus on wall-bounded flows mostly in the turbulent regime. The latter is motivated partially by the scientific interests of the authors, but also by the fact that turbulence near the wall has a particularly rich and complex structure that could potentially be disrupted due to floating point errors. The four use cases cover several important flow classes: internal flows (the channel flow test case), transition to turbulence (the Tollmien-Schlichting wave test case), external flows over bluff bodies (flow around a cylinder), and external flows with influence of compressibility (the NACA-0012 case).
@@
In our analysis, we go below FP32 and also consider precision that is not necessarily implemented in current hardware. To facilitate that, we rely on a software emulation of the precision. However, we have also developed native FP32 versions of Neko, Simson, and, partially, SSDC. This work reveals that naively reducing the precision of all reals may lead to various pitfalls. Therefore, besides presenting simulation results, we also summarize practical experiences in adapting our codes to FP32, in order to aid other members of the community in similar efforts.
@@ 2 Floating-point numbers
In this work, we are interested in the impact of floating point formats and their ability to represent relevant physics in a turbulent flow simulation. A floating-point number is defined by a number of mantissa bits $b$ and exponent bits $b_{e}$ together with one sign bit $s$ dictating the sign of the floating-point number. If we let $e$ be the value of the exponent (as an unsigned $b_{e}$-bit integer), and $c_{i}$ be the $i$th least significant bit of the mantissa, the value for a given normal floating-point number is
@@
We consider floating-point formats between 8 and 64 bits as listed in Table 1. In addition to the normal floating-point numbers, there are also special numbers such as $\pm$ infinity, not a number (NaN) and subnormal numbers.
@@
Floating-point numbers have a great strength in their large range and use of a relative rounding error, compared to an absolute one such as for fixed-point formats. This means that when operations are carried out on numbers of comparable amplitude the effect of rounding error is relatively small. However, issues can, for example, arise when performing computations with two numbers $x$ and $y$ where $x\ll y$. In this case, if $x<\varepsilon y$, the computation may be subject to so-called stagnation. A typical example is when performing summations of long arrays and an individual element in the array is smaller than the total sum, which can yield a significant rounding error. This would, for example, yield inaccurate dot products, large errors after many time steps as well as possibly impact the collection of statistical quantities.
@@
If we consider $\mathbf{u}^{i}$ to be the state of the flow at time step $i$ and time $i\Delta t$. It is a discretized vector of length $n$ $\mathbf{u}\in FP^{n}$ where $FP$ is the set of floating point numbers possible to be represented for a floating-point format FP. The total simulation up to time $i\Delta t$ can then be described as the set of realizations of the flow $\mathbf{U}=\{u^{0},\ldots,u^{i}\}$, where $\mathbf{U}$ is constructed through some map from state $\mathbf{u}^{i}$ to $\mathbf{u}^{i+1}$
@@
In this investigation we make an attempt to see how $f$, which is described by the numerical method, is affected, regardless of and the elements in $\mathbf{u}$ are not the focus, but rather how different terms in the Navier-Stokes equations and the state vector $\mathbf{u}$ are affected by lower floating point precision. This view of a numerical simulation holds for any numerical discretization of the flow. Evaluating different numerical methods such as finite volume, finite elements or similar would in this model equal the choice of the element in the state vector and how the map $f$ is computed.
@@ 3 Methodology
Throughout this work, we consider floating-point precision as a rounding operation from some state $u$ to a rounded state $\tilde{u}$. In our work, for all solvers considered in the following sections, all floating-point numbers below FP32 are emulated with CPFloat [34]. Several opportunities to incorporate rounding into the governing equations exist, and the question is how different terms, such as the advective non-linear term depend on the numerical precision. The nonlinear term is especially important for turbulent cases where $Re$ is large, the equations are chaotic and small disturbances make two trajectories diverge quickly. As such, analytical or deterministic approaches to assess the accuracy of the simulation are no longer applicable. In addition, in a numerical setting where the equations are discretized, the interaction and dependence on an accurate numerical format can be drastic. As such, we consider several different cases to evaluate if a simulation can be run entirely in lower precision, as well as only perturbing specific terms in the governing equations. In particular, we distinguish three approaches:
@@
    State rounding. Casting $\boldsymbol{u}^{i}$ in lower precision, while the solver operates in FP64. Simulating the perturbed system, $\mathbf{u}^{i+1}=f(\tilde{\mathbf{u}^{i}})$, where the state is constrained to a lower floating point precision at each time $i$.
@@
    Term rounding. Different terms in the NS equations, such as the convective or viscous term or both, are represented in lower precision; for example, for the convective term in the incompressible formulation would be computed according to $\widetilde{(\mathbf{v}\cdot\nabla)\mathbf{v}}$ with the rounding operator.
In order to assess how these perturbations impact different numerical schemes and different formulations of the Navier–Stokes equations, we consider four different flow solvers with different discretizations and characteristics. In light of these differences, the specific way the rounding is applied differs somewhat between the codes. These details are discussed for each solver individually.

### v2

@@ Abstract
Modern computing clusters offer specialized hardware for reduced-precision arithmetic, which can significantly speed up the time to solution. This is possible due to a decrease in data movement, as well as the ability to perform arithmetic operations at a faster rate. However, for high-fidelity simulations of turbulence, such as direct and large-eddy simulation, the impact of reduced precision on the computed solution and the resulting uncertainty across flow solvers and different flow cases has not been explored in detail, and limits the optimal utilization of new high-performance computing systems. In this work, the effect of reduced precision is studied using four diverse computational fluid dynamics (CFD) solvers (two incompressible, Neko and Simson, and two compressible, PadeLibs and SSDC) using four test cases: turbulent channel flow at $Re_{\tau}=550$ and higher, forced transition in a channel, flow over a cylinder at $Re_{D}=3900$, and compressible flow over a wing section at $Re_{c}=50000$. We observe that the flow physics are remarkably robust with respect to reductions in lower floating-point precision, and that often other forms of uncertainty, due to, for example, time averaging, often have a much larger impact on the computed result. Our results indicate that different terms in the Navier–Stokes equations can be computed to a lower floating-point accuracy without affecting the results. In particular, standard IEEE single precision can be used effectively for the entirety of the simulation, showing no significant discrepancies from double-precision results across the solvers and cases considered. Potential pitfalls are also discussed.
@@ 1 Introduction
Computational fluid dynamics (CFD) has become an essential tool in both academic research and industry, encompassing a wide range of applications. Over time, a variety of models and numerical methods have been developed to integrate the governing equations of fluid motion, each tailored to specific use cases, desired accuracy, and computational constraints. Yet, until relatively recently, all of them typically relied on IEEE double precision floating point numbers (FP64) to numerically compute the solution.
Due to recent shifts in hardware manufacturing tailored to decrease the energy consumption of floating-point operations and drive up performance for artificial intelligence (AI) applications, hardware support for lower precision floating-point numbers has become increasingly prevalent [1]. On these new platforms, lower precision offers both higher performance and higher energy efficiency, as well as a smaller memory footprint, reducing the amount of data movement necessary. All of these developments ultimately lead to significant savings in both time and energy, and thus provide considerable monetary savings for large-scale computations. As CFD practitioners, where most codes are limited by memory bandwidth due to large relatively sparse linear algebra systems, the use of FP32 instead of FP64 would, for example, move the bandwidth roofline by a factor of 2, indicating an up to 2X of performance would be attainable for a bandwidth-limited code [2]. A simple roofline comparison between two relatively common GPUs, Nvidia A100 and Nvidia GeForce RTX4080, is shown in Figure 1. The simple performance model clearly illustrates how the roofline for the RTX4080, with a significantly lower performance for FP64, improves for FP32 and is comparable to the more expensive A100. Considering that the operational intensity $I$ for a given code also improves with lower precision, this suggests that significant gains can be enabled by lowering precision. In practice, a factor of $2\times$ in performance is rarely achieved due to other factors such as kernel launch latencies or communication overhead in parallel communications.
@@
While there are clear opportunities for performance gains, an important question is to what extent these lower floating point formats can be used for scale-resolving simulations of turbulence without sacrificing the required accuracy. This is what we aim to assess in this work.
@@
This question is currently under active investigation by the research community and can be viewed as part of a broader trend across multiple disciplines [3]. A full review of works on lower-precision arithmetics and CFD is outside the scope of this article. Still, a few selected references are provided for the benefit of the reader. The first example is from simulations of weather and climate. Several articles from Klöwer, Paxton, Düben, and Palmer et al. [4, 5, 6] highlight that lower precision and less exact computing open up for large performance improvements. In the realm of compressible flows, Witherden and Jameson [7] and Vermeire et al. [8] used single precision (FP32) in the context of a high-order flux-reconstruction scheme (discontinuous Galerkin) for implicit large eddy simulation for a wide range of cases using the CFD solver PyFR. Methods that enable mixed-precision computing while maintaining acceptable levels of accuracy have been proposed and investigated, for instance, in the context of the finite volume solver OpenFOAM [9]. Further encouraging results using OpenFOAM were also found by Brogi et al. [10]. Haridas et al. [11] explored the possibility of using neural networks to correct errors introduced using reduced precision arithmetic in simulating fluid dynamic problems. The feasibility of conducting mixed-precision operations in the context of high-order compact finite difference schemes was investigated by Song et al. [12] using the PadeLibs code for CFD. Wang et al. [13] proposed a mixed precision strategy in the finite volume method for unstructured grids that used high precision near solid bodies and lower precision far away from them. Freytag et al. [14] studied the performance and power efficiency of using reduced- and mixed-precision arithmetic for CFD. Bhola and Duraisamy [15, 16] performed analyses of errors incurred due to rounding in mixed precision computations. Walden et al. [17] investigated the speedup achieved on GPUs using the FUN3D CFD code with reduced-precision arithmetic for their memory-bound linear solver kernel. Grout [18] used reduced-precision for the low-order time integration methods employed to construct higher-order methods in the so-called spectral differed correction method used in combustion CFD and studied the resulting rate of convergence. To reduce the communication bottleneck, Reuther et al. [19] in the context of a multiblock RANS solver and Rogowski et al. [20] in the context of discontinuous Galerkin (using the SSDC solver), among others, investigated the impact and performance of reduced precision for parallel communications. More recently, Siklòsi et al. [21] explored the use of mixed precision for compressible turbulent flow computations with an explicit finite difference framework. Finally, for CFD based on the Lattice Boltzmann method rather than solving the Navier-Stokes (NS) equations, [22] demonstrates the possibility to accelerate the solver using lower precision. It is also worth noting that finite volume codes used in industry, both open-source and commercial, offered the possibility to compile in FP32 long before the recent increase in low-precision computer hardware. There are also many mixed-precision algorithms for important linear algebra operations used in CFD, such as preconditioners and iterative solvers, to obtain a result in full double precision [23, 24, 3, 25]. This potentially suggests a wider scale of adoption of FP32 than what is documented in the scientific literature.
@@
Despite these success stories, it is important to acknowledge that, due to the nonlinear nature of the NS equations, cases can be found where numerical precision has a larger-than-expected impact. In particular, in the context of dynamical systems, if there exist several attractors of the flow or a symmetry that is sensitive to small disturbances, lower floating-point precision can be detrimental to the validity of the simulation [26, 27]. However, if the simulation only has one attractor and the results are insensitive to small thermal fluctuations, which, in many ways, can act similarly to low-precision [28], lower precision might be utilized effectively.
@@
The question in the focus of this article is whether reduced precision arithmetic can be successfully employed in direct numerical simulations (DNS) and wall-resolved large-eddy simulation (LES). The authors’ impression is that the DNS community is generally skeptical about abandoning FP64. There are good reasons behind that. DNS are usually conducted to obtain the ground truth for a given flow and study all possible subtleties in its behavior. Therefore, even small errors are considered unacceptable. Furthermore, a new DNS is usually conducted at the limit of the computing budget, and simply trying out a low-precision simulation can be perceived as not worth the risk.
@@
Nevertheless, the main outcome of the DNS is often primarily expressed in a statistical description of the flow, and an argument can be made that it must, to a degree, be robust to small numerical errors, the origin of which may be both discretization and the precision of arithmetics. Indeed, there are already documented successful attempts in using FP32 for DNS of homogeneous isotropic turbulence (HIT) [29, 30, 31, 32], with no discernible differences observed between the FP32 and FP64 simulations. While this is very encouraging, to trigger a shift across the wider high-fidelity CFD community, it is necessary for similar evidence to emerge across a broader range of flow cases and numerical techniques. The goal of the current work is to be a step in that direction and toward a deeper understanding of the potential impact of lower floating precision on high-fidelity CFD. This builds on some preliminary work performed by Karp et al. [33, 34, 35], and should be viewed as an extension of the existing evidence in the literature by investigating a wider range of solvers, flows, and arithmetic precisions.
@@
To that end, our study employs four different solvers (Neko, SSDC, Simson, PadeLibs) which differ in both the formulation of the governing equations and the approach to discretization. We consider four use cases, with a focus on wall-bounded flows mainly in the turbulent regime. The latter is motivated, in part, by the scientific interests of the authors, but also by the fact that turbulence near the wall has a particularly rich and complex structure that could potentially be disrupted due to floating-point errors. The four use cases cover several important flow classes: internal flows (the channel flow test case), transition to turbulence (the Tollmien-Schlichting wave test case), external flows over bluff bodies (flow around a cylinder), and external flows with influence of compressibility (the NACA-0012 case).
@@
To extend the range of arithmetic precisions, we go below FP32 and also consider precision that is not necessarily implemented in current hardware. To facilitate that, we rely on a software emulation of the precision. However, we have also developed native FP32 versions of Neko, Simson, and, partially, SSDC. This work reveals that naively reducing the precision of all reals may lead to various pitfalls. Therefore, in addition to presenting simulation results, we also summarize our practical experiences in adapting our codes to FP32, in order to aid other members of the community in similar efforts. While not used in the work, tools such as the Herbie project [36], which aim at identifying floating-point problems in the code, may be of particular interest when adapting the codes for lower precisions.
@@ 2 Floating-point numbers
In this work, we are interested in the impact of floating-point formats and their ability to represent relevant physics in a turbulent flow simulation. A floating-point number is defined by a number of mantissa bits $b$ and exponent bits $b_{e}$ together with one sign bit $s$ dictating the sign of the floating-point number. If we let $e$ be the value of the exponent (as an unsigned $b_{e}$-bit integer), and $c_{i}$ be the $i$th least significant bit of the mantissa, the value for a given normal floating-point number is
@@
We consider floating-point formats ranging from 8 to 64 bits, as listed in Table 1. In addition to the normal floating-point numbers, there are also special numbers such as $\pm$ infinity, not a number (NaN), and subnormal numbers.
@@
Floating-point numbers have a great strength in their large range and the use of a relative rounding error, compared to an absolute one, such as in fixed-point formats. This means that when operations are carried out on numbers of comparable amplitude, the effect of rounding error is relatively small. However, issues can arise, for example, when performing computations with two numbers $x$ and $y$ where $x\ll y$. In this case, if $x<\varepsilon y$, the computation may be subject to so-called stagnation. A typical example is when performing summations of long arrays, and an individual element in the array is smaller than the total sum, which can yield a significant rounding error. This would, for example, yield inaccurate dot products, large errors after many time steps, as well as possibly impact the collection of statistical quantities.
@@
If we consider $\mathbf{u}^{i}$ to be the state of the flow at time step $i$ and time $i\Delta t$, it is a discretized vector of length $n$, $\mathbf{u}\in FP^{n}$, where $FP$ is the set of floating point numbers possible to be represented for a floating-point format FP. The total simulation up to time $i\Delta t$ can then be described as the set of realizations of the flow $\mathbf{U}=\{u^{0},\ldots,u^{i}\}$, where $\mathbf{U}$ is constructed through some map from state $\mathbf{u}^{i}$ to $\mathbf{u}^{i+1}$
@@
In this investigation we make an attempt to see the impact of floating point precision on different terms in the governing equations and how this affects the final discretized system $f$ and its behavior across numerical methods irrespective of their state representation $\mathbf{u}$. This view of a numerical simulation holds for any numerical discretization of the flow. Evaluating different numerical methods, such as finite volume, finite elements, or similar, would in this model equal the choice of the element in the state vector and how the map $f$ is computed.
@@ 3 Methodology
Throughout this work, we consider floating-point precision as a rounding operation from some state $u$ to a rounded state $\tilde{u}$. In our work, for all solvers considered in the following sections, all floating-point numbers below FP32 are emulated with CPFloat [37]. Several opportunities exist to incorporate rounding into the governing equations, and the question is how different terms, such as the advective nonlinear term, depend on the numerical precision. The nonlinear term is especially important for turbulent cases where $Re$ is large, the solution to the equations is chaotic, and small disturbances cause two trajectories to diverge quickly. As such, analytical or deterministic approaches to assess the accuracy of the simulation are no longer applicable. In addition, in a numerical setting where the equations are discretized, the interaction and dependence on an accurate numerical format can be drastic. As such, we consider several different cases to evaluate whether a simulation can be run entirely in lower precision, as well as whether only specific terms in the governing equations need to be perturbed. In particular, we distinguish three approaches:
@@
[…]
@@
    Term rounding. Different terms in the NS equations, such as the convective or viscous term or both, are represented in lower precision; for example, the convective term in the incompressible formulation would be computed according to $\widetilde{(\mathbf{v}\cdot\nabla)\mathbf{v}}$ with the rounding operator.
To assess the impact of these perturbations on various numerical schemes and formulations of the incompressible and compressible Navier–Stokes equations, we consider four different flow solvers with distinct discretizations and characteristics. In light of these differences, the specific way the rounding is applied differs between the codes. These details are discussed for each solver individually.